Is it possible to have a Binomial Distribution in the following cases:


i. Mean = 5 and Variance = 2.5


ii. P(X = – 2) = 0.1467


iii. P(X less than or not equal to 1) = 0.18


iv. Three parameters i.e. n, p and q


v. P(X = 1) = 0.35


vi. P(X Greater than or equal to 1) = 1.27


vii. .P(X = 3.5) = 0.19


viii. P(1   less than or not equal to   X     less than or not equal to    4) = 0.79.


1
Expert's answer
2022-05-24T15:29:05-0400

i.


"\mu=np=5, Var(X)=npq=2.5"

"q=0.5, 0<0.5<1"

It is possible.


ii. "X" must be nonnegative.

It is impossible.


iii.


"P(X<1\ or X\not=1)=1-P(X=1)=0.18"

It is possible.


iv. Three parameters "n, p, q" can define a Binomial Distribution.

It is possible.


v. "P(X=1)=0.35"

It is possible.


vi. "P(X\ge1)=1.27"


"0\le p\le 1"

It is impossible.


vii. "P(X=3.5)=0.19"

"X" must be integer.

 It is impossible.


viii.


"P((X<1\ or X\not=1)\cap(X<4\ or X\not=4))"

"=1-P(X=1)-P(X=4)=0.79"

It is possible.